OpenStudy (anonymous):

evauate the expression 3+6x(-5)^2

OpenStudy (anonymous):

can someone please help

OpenStudy (unklerhaukus):

is that an 'x' or a 'times'?

OpenStudy (unklerhaukus):

i think its a times,

OpenStudy (unklerhaukus):

you can simplify this term first (-5)^2

OpenStudy (anonymous):

so wat is the answer

OpenStudy (anonymous):

dude, is it a variable or a multiplication symbol?

OpenStudy (anonymous):

I think it can : 3+6x*(-5)^2=3+6x*25=3+175x Or if have : 3+6x^(-5)^2=3+6x^25 But I think its the first ;) Thank you guys ;)